the base of a triangle is 6cm greater than the height. the area is 20cm. find the height and the length of the base
Answer by checkley77(12569)

area=bh/2
20=(h+6)h/2
20=(h^2+6h)/2 cross multiply.
h^2+6h=2*20
h^2+6h=40
h^2+6h-40=0
(h+10)(h-4)=0
h-4=0
h=4 ans. for the height.
4+6=10 ans. for the base.
Proof:
20=4*10/2
20=40/2
20=20